OpenStudy (anonymous):

Find a point on the curve defined by function y = x^2 that has the minimum distance to the line defined by y = 2x - 4. (c) Solve this problem using Lagrange multiplier approach.

OpenStudy (anonymous):

When I solved this problem, I ended up with (1, -2), but I think the answer should be (1,1).

OpenStudy (ybarrap):

Lagrange Equation: $$ \Lambda(x,y,\lambda) = f(x,y) + \lambda \cdot \Big(g(x,y)-c\Big) $$ Let \(f(x,y)=\sqrt{\left (x^2\right )^2+(2x-4)^2}\), which is the distance from \(x^2\) to \(2x-4\). We want to minimize this function. Our constraint is $$ y=x^2\\ $$ Or $$ y-x^2=0 $$ Therefore, $$ \Lambda(x,y,\lambda)=\sqrt{\left (x^2\right )^2+(2x-4)^2}+\lambda \cdot \Big(y-x^2)\Big) $$ Now let's compute partials: $$ \Lambda_x={1\over2\sqrt{\Big(x^4+(2x-4)^2\Big )}}\Big(4x^3+8x-16\Big)-2x\lambda\\ \Lambda_y=\lambda\\ \Lambda_\lambda=y-x^2 $$ Setting each equal to zero (you'll need to work out these details): $$ \Lambda_x=0\\ \Lambda_y=0\\ \Lambda_\lambda=0\\ \implies x\approx1.1795 $$ Hope this helps.